# Feature-based One-For-All: Heterogeneous Distillation Across Vision Architectures

### v1

@@ Introduction
On CIFAR-100, we compare our FOFA method with eight other KD methods, encompassing a total of twelve heterogeneous teacher-student pairings. As shown in Table 1, FOFA outperforms other methods on most teacher-student pairs, resulting in an average improvement of 8.17%. For instance, we achieve an accuracy of 83.50% on the ConvNeXt-T - ResMLP-S12 pair and 79.59% on the ConvNeXt-T - DeiT-T pair, surpassing not only the other feature-based methods but also obtaining an improvement of 2.28% and 3.83% compared to the prior SOTA logits-based method OFA. Specifically, FOFA does not achieve as much improvement on CNN-based students as it does on ViT-based and MLP-based students. Unlike the other two types of students, CNN-based students need additional patch-embedding layers, we speculate that these additional parameters could result in prolonged training epochs, thus only having a small improvement in current settings.
@@
Due to computational limitations, we select four teacher-student settings to experiment with on ImageNet. The results are presented in Table 2. With an increase in the volume of training data, the majority of methods exhibit a favorable average enhancement. Among them, FOFA attains SOTA performance across most teacher-student pairs. For instance, we achieve 72.22% on MobileNetV2 and 77.59% on ResMLP-S12 when distilling from the teacher model Mixer-B/16 and Swin-T, which is 3.35% and 0.94% higher than the baselines, proving our generalizability on a larger dataset.
@@
Results on Object Detection
We extend our method to object detection and follow the setting of DKD (Zhao et al. 2022) to perform distillation between the teacher’s and student’s backbone features. The results are presented in Table 3. Logits-based methods, as mentioned in DKD, can only achieve inferior performance as they disregard the intermediate features. For example, OFA can not reach good performance owing to the excessive focus on semantic details instead of spatial aspects, which is crucial for detection. Conversely, feature-based methods demonstrate superior performance, with our approach achieving SOTA results. We observe an enhancement of 2.36% for ResNet18 and 3.50% for MobileNetV2 when distilling from the Swin-T teacher model. This underscores that by mitigating view mismatch and teacher unawareness issues, the feature-mimicking technique can effectively leverage the abundant intermediate features for improved performance across classification and downstream tasks.
@@
Table 4: The effectiveness of our proposed RAA, AFP modules on CIFAR-100 with ConvNeXt-T - DeiT-T as the teacher-student pair.
Ablation Study
Effectiveness of each proposed module
@@
Effectiveness of number and position of AFP
@@
In RAA, every stage feature of the student model is projected into a predefined sequential form with shape $\mathbb{R}^{\frac{N_{q}}{4}\times d}$ to perform the subsequent attention calculation. As the value of $N_{q}$ increases, the feature can be further subdivided, offering more detailed insights into the original feature. However, this also leads to a larger attention matrix for computation. Here, we investigate the influence on performance and memory usage of the number $N_{q}$ on RAA by varying the number from 36 to 144. As shown in Table 6, the performance consistently improves as the $N_{q}$ increases, alongside a proportional increase in memory consumption, and we can already achieve SOTA on CIFAR-100 when $N_{q}$ is 64 with the least memory usage. In default, we set $N_{q}$ as 64 on CIFAR-100 and set it to a higher value 144 for the larger and more complex datasets like ImageNet and COCO.
Visualization of attention map within RAA
@@
Performing feature matching directly in heterogeneous KD is inefficient, as distinct architectures exhibit varying inductive biases that result in different feature perspectives. In FOFA, the student model is enabled to learn how to blend the features of different regions from different stages via the attention mechanism to get a similar perspective to the teacher model. The attention matrices within the RAA of distinct teacher-student model pairs are illustrated in Fig 3. It can be noted that various teacher-student model combinations lead to diverse patterns of attention. For example, the MLP-based student ResMLP-S12 focuses on the consistent spatial position across stages when learning from the CNN-based teacher ConvNeXt-T, resulting in numerous diagonal lines, indicating that ResMLP-S12 learns to aggregate local information to achieve a similar perspective as ConvNeXt-T. Conversely, the CNN-based student ResNet18 tends to aggregate additional features from neighboring patches within the stage for gathering more global information when learning from the ViT-based teacher Swin-T, producing grid-like formations in the attention map. The ViT-based student DeiT-T also demonstrates a similar tendency to CNN-based students when being instructed by an MLP-based teacher Mixer-B16, however, DeiT-T gathers more information from stage 3 on most of the queries. This might indicate most information is stored in stage 3.
Conclusion
In this paper, we propose a feature-based one-for-all (FOFA) framework to facilitate feature-level supervision in the context of heterogeneous KD. First, we introduce region-aware attention to blend the stage features of the student model to align with those of the teacher model, thereby alleviating the issue of view mismatch prevalent between the two models. Moreover, we incorporate the prompt tuning technique to address the teacher unawareness problem and enable to adapt the teacher model’s feature utilizing feedback from the student model with minimal parameters. Through the integration of these two components, we can realize feature-level distillation in heterogeneous KD scenarios and achieve promising performance in classification and downstream tasks.

### v3

@@ 4.1 Results on Image Classification
On CIFAR-100, we compare our PAT method with eight other KD methods, encompassing a total of twelve heterogeneous teacher-student pairings. As shown in Table 1, PAT outperforms other methods on most teacher-student pairs, resulting in an average improvement of 8.17%. For instance, we achieve an accuracy of 83.50% on the ConvNeXt-T - ResMLP-S12 pair and 79.59% on the ConvNeXt-T - DeiT-T pair, surpassing not only the other feature-based methods but also obtaining an improvement of 2.28% and 3.83% compared to the prior SOTA logits-based method OFA. Specifically, PAT does not achieve as much improvement on CNN-based students as it does on ViT-based and MLP-based students. Unlike the other two types of students, CNN-based students need additional patch-embedding layers, we speculate that these additional parameters could result in prolonged training epochs, thus only having a small improvement in current settings.
@@
Due to computational limitations, we select four teacher-student settings to experiment with on ImageNet. The results are presented in Table 2. With an increase in the volume of training data, the majority of methods exhibit a favorable average enhancement. Among them, PAT attains SOTA performance across most teacher-student pairs. For instance, we achieve 72.22% on MobileNetV2 and 77.59% on ResMLP-S12 when distilling from the teacher model Mixer-B/16 and Swin-T, which is 3.35% and 0.94% higher than the baselines, proving our generalizability on a larger dataset.
@@
4.2 Results on Object Detection
We extend our method to object detection and follow the setting of DKD [37] to perform distillation between the teacher’s and student’s backbone features. The results are presented in Table 3. Logits-based methods, as mentioned in DKD, can only achieve inferior performance as they disregard the intermediate features. For example, OFA can not reach good performance owing to the excessive focus on semantic details instead of spatial aspects, which is crucial for detection. Conversely, feature-based methods demonstrate superior performance, with our approach achieving SOTA results. We observe an enhancement of 2.36% for ResNet18 and 3.50% for MobileNetV2 when distilling from the Swin-T teacher model. This underscores that by mitigating view mismatch and teacher unawareness issues, the feature-mimicking technique can effectively leverage the abundant intermediate features for improved performance across classification and downstream tasks.
@@ 4.2 Results on Object Detection
Table 4: The effectiveness of our proposed RAA, AFP on CIFAR-100 with ConvNeXt-T - DeiT-T as the teacher-student pair.
5 Ablation Studies
5.1 Effectiveness of each proposed module
@@
5.2 Effectiveness of number and position of AFP
@@ 5.3 Number of queries on RAA
In RAA, each stage feature of the student is projected into a predefined sequence of shape $\mathbb{R}^{\frac{N_{q}}{4}\times d}$ for attention computation. Increasing $N_{q}$ allows finer subdivision of features, offering more detailed representations, but also leads to larger attention matrices and higher computational cost. We evaluate the impact of varying $N_{q}$ from 36 to 144 on performance and memory usage. As shown in Table 6, performance improves steadily with larger $N_{q}$, accompanied by proportional memory growth. Notably, setting $N_{q}=64$ achieves SOTA on CIFAR-100 with minimal memory usage. By default, we use $N_{q}=64$ for CIFAR-100 and $N_{q}=144$ for larger datasets such as ImageNet and COCO.
5.4 The performance of PAT without KL loss
Teacher Swin-T ConvNeXt-T ConvNeXt-T
Student ResNet18 DeiT-T ResMLP-S12
KD 78.74 72.99 72.25
FitNet 78.87 60.78 45.47
PAT w/o KL Loss 79.14 75.98 81.79
PAT w/ KL Loss 81.22 79.10 83.50
Table 7: The effectiveness of loss $L_{KL}$ on PAT on CIFAR-100.
PAT is not purely feature-based, as it also incorporates a logits-based KL-divergence loss, as shown in Eq.8. This raises the question of whether performance gains stem primarily from the logits-matching loss rather than the feature-level alignment. To investigate, we train PAT without the KL loss. As shown in Table7, adding the KL loss improves performance. However, even without it, PAT still achieves performance close to SOTA. For best results, we include the KL loss by default to further enhance overall effectiveness.
5.5 Complexity Analysis
Method Extra Para. (M) Epoch (s) Acc.
KD 0 65.75 72.99
OFA 3.78 76.16 75.76
FitNet 1.11 90.45 60.78
PAT 14.48 208.57 79.59
Table 8: The complexity of KD methods with ConvNeXt-T - DeiT-T as the teacher-student pair on CIFAR-100.
We also explore the complexity of our method. As shown in Table 8, it is found that logits-based methods, which rely solely on logits for distillation, require significantly fewer parameters compared to feature-based methods, leading to faster training times. In contrast, feature-based methods involve more parameters, which extends the training duration. Although PAT consumes the most parameters, it also achieves the highest performance. Notably, these additional parameters only appear during training and do not add any overhead to the student model during inference.
For the trade-off between the complexity and performance of PAT, as mentioned in section  5.2 and  5.3, we can achieve faster training by lowering the number of AFP modules, or by decreasing the queries number on RAA.
5.6 Visualization of attention map within RAA
@@
Direct feature matching in heterogeneous KD is suboptimal due to differing inductive biases across architectures, leading to mismatched feature representations. To address this, PAT leverages attention to guide the student in blending features from different regions and stages, aligning its perspective with the teacher. As shown in Fig. 3, attention patterns vary across teacher-student pairs. For example, the MLP-based student ResMLP-S12 learns from the CNN-based teacher ConvNeXt-T by attending to consistent spatial positions across stages, forming diagonal patterns that reflect local aggregation. In contrast, the CNN-based student ResNet18, learning from the ViT-based teacher Swin-T, captures features from neighboring patches, resulting in grid-like patterns indicative of global context aggregation. Similarly, the ViT-based student DeiT-T, when guided by the MLP-based teacher Mixer-B16, exhibits global attention with a strong focus on stage 3, suggesting that key information is concentrated there.
6 Conclusion
In this paper, we propose a perspective-aware teaching (PAT) framework to facilitate feature-level supervision in the context of heterogeneous KD. First, we introduce region-aware attention to blend the stage features of the student model to align with those of the teacher model, thereby alleviating the issue of view mismatch prevalent between the two models. Moreover, we incorporate the prompt tuning technique to address the teacher unawareness problem and enable to adapt the teacher model’s feature utilizing feedback from the student model with minimal parameters. Through the integration of these two components, we can realize feature-level distillation in heterogeneous KD scenarios and achieve promising performance in classification and downstream tasks.
7 Acknowledgments
This work is partially supported by the National Science and Technology Council, Taiwan, under Grant: NSTC-114-2640-E-A49-011, NSTC-114-2218-E-A49-017, and NSTC-112-2628-E-002-033-MY4, and was financially supported in part by the Center of Data Intelligence: Technologies, Applications, and Systems, National Taiwan University (Grants: 114L900901/114L900902/114L900903), from the Featured Areas Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education, Taiwan.